OpenStudy (anonymous):

Karim wants to buy a shirt that normally sells for $46.50. It is on sale for 1/3 off. How much does Karim need to pay to buy the shirt on sale? $15.50 $22.50 $31.00 $41.50

OpenStudy (swolesammy):

C

OpenStudy (mathstudent55):

First, find 1/3 of $46.50. Then subtract that amount from $46.50.

OpenStudy (anonymous):

idk the formula

OpenStudy (texaschic101):

1/3 of 46.50....." of " means multiply 1/3 * 46.50 = 46.50/3 = 15.50.....that is how much the sale is....15.50 off of 46.50. 46.50 - 15.50 = 31.00.....so with the discount applied, Karim would have to pay $31.00. questions ?

OpenStudy (anonymous):

Its the Price of the shirt x 1/3 46.50 x 1/3 = 15.5 Price he needs to pay = $(46.50 - 15.5) = $31.50. So its option C.

OpenStudy (anonymous):

No Ma'am no questions im doing a topic test review and i hate math idk the the formulas it so confusing